The ratio of magnetic field due to a smaller bar magnet in the end on position to broad side on position is ...A)1/4. B) 1/2 c) 1. D) 2

Grade:12th pass

3 Answers

ROSHAN MUJEEB
3 years ago
2
SACHIN KUMAR MISHRA
2 years ago
End-on position means axial position while broad - side means equilateral position from a bar magnet.
Let the length of a bar magnet is l. Hence magnetic field on the axis of bar magnet at a distance r from the centre of magnet where  (l
 
B_1 = \frac{\mu _0}{4\pi }\frac{2M}{r^3}          …......(1)
The magnetic field on the equilateral axis at a distance r from centre of a bar magnet 
B_2 = \frac{\mu _0}{4\pi }\frac{M}{r^3} \ \ \ \ \ \ .........(2)
Dividing equation (1) and (2)
\frac{B_1}{B_2} \ = \frac{2}{1}
 
We know that magnetic field on the axis of a bar magnet at a distance 
SACHIN KUMAR MISHRA
2 years ago
End-on position means axial position while broad - side means equilateral position from a bar magnet.
Let the length of a bar magnet is l. Hence magnetic field on the axis of bar magnet at a distance r from the centre of magnet where  (l
 
B_1 = \frac{\mu _0}{4\pi }\frac{2M}{r^3}          ...…......(1)
The magnetic field on the equilateral axis at a distance r from centre of a bar magnet 
B_2 = \frac{\mu _0}{4\pi }\frac{M}{r^3} \ \ \ \ \ \ .........(2)
Dividing equation (1) and (2)
\frac{B_1}{B_2} \ = \frac{2}{1}